the circumference of a circle is 14.5cm
a) calculate the diameter of the circle
b) calculate the radius of the circle
c) calculate the area of the cirle

Answer by user_dude2008(1825)

a) calculate the diameter of the circle


C=2*pi*r

14.5=2*3.14*r

14.5/(2*3.14)=r

14.5/(6.28)=r

2.31=r

r=2.31

diameter: d=2*r=2(2.31)=4.62

-------------------

b) from part a) radius: r=2.31

-------------------
c)
Area:
A=pi*r^2

A=(3.14)(2.31)^2

A=(3.14)(5.3361)

A=16.755